A large cylindrical tank contains 0.760 m3 of nitrogen gas at 30.0 ∘C and 1.50×105 Pa (absolute pressure). The tank has a tight-fitting piston that allows the volume to be changed.

What will be the pressure if the volume is decreased to 0.470 m3 and the temperature is increased to 151 ∘C?
Answer #1

Solution) V1 = 0.760 m^3

T1 = 30°C = 30+273 = 303 K

P1 = 1.5×10^(5) Pa

V2 = 0.470 m^3

T2 = 151°C = 151+273 = 424 K

P2 = ?

From

PV = nRT

As n and R are constant

((P1)(V1))/(T1) = ((P2)(V2))/(T2)

(1.5×10^(5)×0.760)/(303) = (P2)(0.470/424)

376.24 = 0.0011(P2)

P2 = (376.24/0.0011)

P2 = 3.42×10^(5) Pa